# Improving Low-Resource Sequence Labeling with Knowledge Fusion and Contextual Label Explanations

### v2

@@ 5 Conclusion
In this paper, we propose a novel framework that integrates an LLM-based knowledge enhancement workflow with a span-based sequence labeling model. Our approach improves model performance by generating contextual interpretations of target entities and annotating extension labels. Additionally, our KnowFREE model effectively incorporates extension label features to enhance extraction capabilities. Extensive experiments on Chinese few-shot sequence labeling datasets demonstrate that our method achieves state-of-the-art performance, showcasing its effectiveness and efficiency.
@@ Appendix A Sequence Labeling with LLMs
Dataset / Model Weibo Youku Taobao Resume CMeEE UD PKU MSR
GLM3-6B 3.28 9.84 9.21 13.71 2.41 1.15 3.76 6.69
GLM4-9B 12.43 31.69 14.90 47.12 16.28 26.38 26.18 31.12
Qwen2.5-14B 31.06 51.00 5.89 51.67 22.38 10.06 2.33 3.02
@@
Table 3: Results on sequence labeling tasks with vanilla LLMs.
Dataset / Model Weibo Youku Taobao Resume
$\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000
GLM3-6B (LoRA) 12.29 20.44 28.22 34.52 38.62 39.29 29.91 30.02 30.03 39.92 40.76 39.98
GLM4-9B (LoRA) 49.20 54.54 60.30 73.79 74.56 74.84 58.84 63.63 69.20 83.31 86.24 88.27
Qwen2.5-14B (LoRA) 50.93 56.87 56.30 72.89 75.10 78.83 61.58 59.43 66.98 84.45 88.12 88.65
Dataset / Model CMeEE-v2 PKU MSR UD
$\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000
GLM3-6B (LoRA) 12.75 13.30 13.55 20.62 26.27 29.21 23.53 26.04 29.38 32.20 37.57 37.81
GLM4-9B (LoRA) 50.67 54.50 56.22 71.41 73.68 75.02 75.70 78.64 80.94 72.32 77.50 79.69
Qwen2.5-14B (LoRA) 53.59 53.80 59.31 71.94 72.71 75.91 77.83 79.89 80.07 74.40 77.04 80.70
@@
Table 4: Performance of LLM fine-tuning with LoRA on sequence labeling tasks.
@@
To evaluate the performance of directly using LLMs for sequence labeling tasks in low-resource scenarios, we present the results of vanilla LLMs in Table 3 and the performance of LoRA-finetuned models trained on sampled datasets in Table 6. During zero-shot inference, LLMs extract entities based on the input text and target labels. As shown in Table 3, vanilla LLMs exhibit significantly poor performance on all datasets, likely due to their limited understanding of target label definitions. Therefore, for domain-specific sequence labeling tasks, incorporating few-shot examples into the prompt or fine-tuning the model with LoRA could be a more effective and practical approach.
@@
In the experiments with LoRA fine-tuning, all LLMs show performance improvements compared to zero-shot inference. Compared to other LLMs, GLM3-6B still underperforms in LoRA fine-tuning, likely due to its weaker ability to align with the target domain of sequence labeling. As a result, the synthesized data produced by GLM3-6B contains a considerable amount of irrelevant information. With $\mathcal{K}$=250, Qwen2.5-14B demonstrates outperforms other LLMs across all datasets, suggesting that models with larger parameters tend to show enhanced initial performance in low-resource contexts. Nonetheless, with the sample size increases, the performance improvements of Qwen2.5-14B on datasets such as Weibo, Taobao, and MSR fell short compared to GLM4-9B. This could be due to variations in model performance when applied to different domain-specific data distributions. It is important to highlight that the performance of LLMs on most datasets was still below that of conventional sequence labeling baselines, including the the relatively simple BERT-CRF. The findings from the main results indicate that the knowledge contained in LLMs can significantly improve sequence labeling performance. However, they lack the necessary expertise and alignment capabilities for handling domain-specific datasets. Due to biases in domain data distribution, LLMs struggle to identify target entities in particular fields as efficiently as conventional sequence labeling techniques.
@@
Appendix B Visual Analysis of Enriched Explanation Synthesis
Figure 5: t-SNE visualization of the training, test and enriched explanation samples under different sampling sizes. The synthetic enriched explanation samples are generated by GLM3-6B, and they are represented by the “Synthetic” in the legend.
In this section, we use the BERT-based embedding model, text2vec (Ming, 2022), to generate sentence embeddings for the original training and test samples, as well as the enriched explanation samples from the Weibo, Youku, Taobao, Resume, and CMeEE-v2 datasets. These embeddings are projected into a two-dimensional space using the t-SNE algorithm, with the results shown in Figure 5. At $\mathcal{K}$=250, the sparse distribution of training samples in datasets such as Weibo, Youku, Taobao, and Resume fails to fully cover the semantic space of the test sets, the limitation is observed across all datasets. However, the synthesized samples effectively bridge these gaps in the semantic space, leading to significant performance improvements in low-resource scenarios. As $\mathcal{K}$ increases, the training samples begin to provide more comprehensive coverage of the semantic space for most datasets. At $\mathcal{K}$=1000, however, semantic distribution discrepancies are noticeable between some training and synthesized samples in datasets like Youku, Taobao, and CMeEE-v2, potentially introducing noise that may adversely affect model performance. On the Weibo dataset, certain regions of the test set’s semantic space remain underrepresented by the original training samples, even at $\mathcal{K}$=1000. This explains why models trained with synthesized samples continue to exhibit notable performance advantages over those trained without synthesized samples at this sample size.
These observations further underscore the effectiveness of the enriched explanation synthesis method in improving model performance under low-resource conditions. However, they also highlight that as the sample size increases, the potential adverse effects of synthesized samples, such as semantic noise, become more pronounced.
Appendix C Impact of the Number of Heads in Local Attention
@@
Figure 7: Performance variation with different numbers of heads in local attention module.
To analyze the impact of different numbers of attention heads on model performance, we conducted experiments on four flat NER datasets: Weibo, Youku, Taobao, and Resume, using a sampling size of 1000. The model was trained on the sampled data with extension labels extracted by GLM3-6B, and the results are presented in Figure 7.
The results show that the number of attention heads has a moderate effect on performance. Overall, as the number of heads increases, the model’s performance improves slightly, though the gains are not particularly significant. Additionally, it is important to consider how feature vectors are distributed across the heads. Based on these findings, we chose to use ten attention heads in this study.
Appendix D Comparison of different data synthesis strategies
To evaluate the effectiveness of the enriched explanation synthesis strategy, we reproduced and compared two LLM-based data synthesis methods designed for sequence labeling tasks: LLM-DA (Ye et al., 2024) and ProgGen (Heng et al., 2024). These methods were used to synthesize data from k-shot samples of the original datasets. For consistency, all synthesized samples were annotated using the KnowFREE model trained on the original data without synthesized samples. We conducted experiments on the Weibo, Youku, Taobao, and Resume datasets, all results are presented in Figure 6. The results show that the performance gains from all data synthesis strategies decrease as the number of samples increases. Notably, in scenarios with $\text{k}\leq 30$ on the Youku and Taobao datasets, both LLM-DA and ProgGen lead to performance degradation compared to models trained without synthesized data. This suggests that the synthesized samples generated by these methods may contain inherent semantic distribution biases, which diminish their effectiveness in enhancing performance in certain low-resource domains.
@@ Appendix D Comparison of different data synthesis strategies
Appendix E Visualization of the Logits with Extension Labels
@@
Appendix F Statistics of Datasets
@@
Appendix G More Experiment Settings
In this section, we describe the additional experimental parameter settings for our method. In the sequence labeling model, we distinguish between the learning rate for the PLM and other modules, setting them to 2e-5 and 1e-3, respectively. For the weight $\alpha$ of extension labels, we introduce a dynamic weight calculation mechanism to handle the influence of frequently occurring extension labels (e.g., POS tags). These frequent labels can affect the gradient calculation, leading to reduced attention to target labels. To address this, we calculate the count $C_{i}$ of each extension label and the average count $\hat{C}$ of target labels, and then compute the weight coefficient $\alpha_{i}$ as follows:

### v4

@@ 5 Conclusion
In this paper, we propose a novel framework that integrates an LLM-based knowledge enhancement workflow with a span-based sequence labeling model. Our approach improves model performance by generating contextual interpretations of target entities and annotating extension labels. Additionally, our KnowFREE model effectively incorporates extension label features to enhance extraction capabilities. Extensive experiments demonstrate that our method achieves state-of-the-art performance, showcasing its effectiveness and efficiency.
@@
Acknowledgments
This work is supported by National Natural Science Foundation of China (U23B2048, 62402011) and High-performance Computing Platform of Peking University. Bin Cui is the corresponding author.
@@ Appendix A Sequence Labeling with LLMs
Model Weibo Youku Taobao Resume CMeEE UD PKU MSR
ChatGLM3-6B 3.28 9.84 9.21 13.71 2.41 1.15 3.76 6.69
Llama3-8B-Instruct 15.06 13.89 9.67 41.45 11.27 14.06 23.88 28.84
GLM4-9B-Chat 12.43 31.69 14.90 47.12 16.28 26.38 26.18 31.12
Qwen2.5-14B-Instruct 31.06 51.00 5.89 51.67 22.38 10.06 2.33 3.02
Llama3.1-70B-Instruct 34.69 49.54 13.77 66.60 40.86 37.82 3.96 7.44
Deepseek-V3 33.33 10.91 15.91 60.88 42.20 38.66 5.96 2.45
@@
Table 4: Results on sequence labeling tasks with vanilla LLMs on zero-shot learning.
Model Weibo Youku Taobao Resume
250 500 1000 250 500 1000 250 500 1000 250 500 1000
ChatGLM3-6B (LoRA) 12.29 20.44 28.22 34.52 38.62 39.29 29.91 30.02 30.03 39.92 40.76 39.98
GLM4-9B-Chat (LoRA) 49.20 54.54 60.30 73.79 74.56 74.84 58.84 63.63 69.20 83.31 86.24 88.27
Llama3-8B-Instruct (LoRA) 10.09 17.08 44.24 49.59 60.97 64.33 23.84 34.65 42.98 65.36 70.82 74.86
Qwen2.5-14B-Instruct (LoRA) 50.93 56.87 56.30 72.89 75.10 78.83 61.58 59.43 66.98 84.45 88.12 88.65
Model CMeEE-v2 PKU MSR UD
250 500 1000 250 500 1000 250 500 1000 250 500 1000
ChatGLM3-6B (LoRA) 12.75 13.30 13.55 20.62 26.27 29.21 23.53 26.04 29.38 32.20 37.57 37.81
GLM4-9B-Chat (LoRA) 50.67 54.50 56.22 71.41 73.68 75.02 75.70 78.64 80.94 72.32 77.50 79.69
Llama3-8B-Instruct (LoRA) 23.27 41.52 44.41 62.50 64.53 68.88 66.17 68.52 69.72 32.16 35.51 44.85
Qwen2.5-14B-Instruct (LoRA) 53.59 53.80 59.31 71.94 72.71 75.91 77.83 79.89 80.07 74.40 77.04 80.70
@@
Table 5: Performance of LLM fine-tuning with LoRA on sequence labeling tasks with source datasets.
Model Weibo Youku Taobao Resume
250 500 1000 250 500 1000 250 500 1000 250 500 1000
ChatGLM3-6B (LoRA) 5.81 6.11 6.09 11.98 12.33 12.37 3.69 3.92 3.98 10.88 12.69 13.98
GLM4-9B-Chat (LoRA) 40.22 58.18 62.38 69.81 75.72 74.68 59.03 63.22 69.03 83.47 86.34 88.10
Llama3-8B-Instruct (LoRA) 22.08 12.42 11.79 18.97 19.85 20.35 20.22 16.76 16.12 49.10 46.11 45.65
Qwen2.5-14B-Instruct (LoRA) 46.22 51.20 56.06 31.92 19.40 20.60 62.25 62.26 65.10 82.50 87.28 88.54
Model CMeEE-v2 PKU MSR UD
250 500 1000 250 500 1000 250 500 1000 250 500 1000
ChatGLM3-6B (LoRA) 1.14 1.02 1.16 1.89 1.78 1.69 4.71 7.62 7.75 5.65 6.17 6.68
GLM4-9B-Chat (LoRA) 50.02 54.42 56.79 70.57 73.18 75.01 75.33 78.28 80.26 70.94 78.41 79.08
Llama3-8B-Instruct (LoRA) 20.69 19.86 19.29 10.56 9.48 8.69 27.36 26.37 25.64 31.25 30.18 30.02
Qwen2.5-14B-Instruct (LoRA) 52.17 54.19 60.83 72.01 72.81 75.55 77.48 79.22 80.95 74.09 77.41 80.55
Table 6: Performance of LLM fine-tuning with LoRA on sequence labeling tasks augmented with synthetic datasets.
@@
To evaluate the performance of directly using LLMs for sequence labeling tasks in low-resource scenarios, we present the results of vanilla LLMs in Table 4 and the performance of LoRA-finetuned models trained on sampled datasets in Table 5. During zero-shot inference, LLMs extract entities based on the input text and target labels. As shown in Table 4, vanilla LLMs exhibit significantly poor performance on all datasets, likely due to their limited understanding of target label definitions. Moreover, models with different parameter scales show varying performance across datasets, and no clear positive correlation is observed between model size and performance. This may be improved through more advanced designs of the sequence labeling prompts used in our setting. Therefore, for domain-specific sequence labeling tasks, incorporating few-shot examples into the prompt or fine-tuning the model with LoRA could be a more effective and practical approach.
@@
In the experiments with LoRA fine-tuning, all LLMs show performance improvements compared to zero-shot inference. Compared to other LLMs, ChatGLM3-6B still underperforms in LoRA fine-tuning, likely due to its weaker ability to align with the target domain of sequence labeling. As a result, the synthesized data produced by ChatGLM3-6B contains a considerable amount of irrelevant information. With the 250-sample setting, Qwen2.5-14B-Instruct demonstrates outperforms other LLMs on most datasets, suggesting that models with larger parameters tend to show enhanced initial performance in low-resource contexts. Nonetheless, with the sample size increases, the performance improvements of Qwen2.5-14B-Instruct on datasets such as Weibo, Taobao, and MSR fell short compared to GLM4-9B-Chat. This could be due to variations in model performance when applied to different domain-specific data distributions. It is important to highlight that the performance of LLMs on most datasets was still below that of conventional sequence labeling baselines, including the the relatively simple BERT-CRF. The findings from the main results indicate that the knowledge contained in LLMs can significantly improve sequence labeling performance. However, they lack the necessary expertise and alignment capabilities for handling domain-specific datasets. Due to biases in domain data distribution, LLMs struggle to identify target entities in particular fields as efficiently as conventional sequence labeling techniques.
In addition, we perform LoRA-based SFT on both source and synthetic data, with the results presented in Table 6. For GLM4-9B-Chat and Qwen2.5-14B-Instruct, fine-tuning on synthetic data does not consistently outperform the results reported in Table 5. This suggests that performance is influenced more by training dynamics and stochastic variations during generation and fine-tuning than by the presence of additional synthetic data. Furthermore, we observe a notable performance decline in ChatGLM3-6B and Llama3-8B-Instruct. Prediction analysis reveals that these models tend to overpredict non-target entities, likely because the frequent co-occurrence of entities in both synthetic and original samples makes them overly sensitive to spurious patterns.
@@
Appendix B Results on Full Datasets
Model Weibo Youku Taobao Resume CMeEE PKU MSR UD CoNLL’03 MIT-Movie
W2NER 72.59 83.62 88.27 96.88 72.97 95.51 97.72 95.00 91.71 74.62
CNN Nested NER 72.31 83.79 88.86 96.67 73.83 93.75 97.69 94.88 91.16 74.86
DiFiNet 73.33 83.69 88.19 96.59 72.29 94.86 97.28 94.85 90.72 74.49
KnowFREE 73.87 84.52 88.97 96.82 73.92 96.59 97.72 95.93 92.28 75.32
KnowFREE-F (Deepseek-V3) 74.15 84.57 89.12 96.93 73.95 96.67 97.76 96.02 92.27 75.38
Table 7: Results of sequence labeling tasks on full datasets. The bold values indicate the best performance.
To evaluate the effectiveness of Label Extension Annotation at the full data scale, we conducted additional experiments comparing our method with W2NER, CNN Nested NER, and DiFiNet. The results are shown in Table 7. While performance improvements become less pronounced on certain datasets (e.g., Weibo, Youku, and Resume) compared to the 1000-sample setting, KnowFREE and KnowFREE-F (Deepseek-V3) still outperform all baseline methods on the full datasets, demonstrating their robustness even under high-resource conditions. Although the impact of label extension annotation diminishes as the dataset size increases, it consistently offers improvements over the vanilla KnowFREE, confirming its continued utility. As for enriched explanation synthesis, one of our main motivations was to investigate its performance boundary in low-resource settings. Our analysis shows that its benefits significantly decrease as the sample size grows, with little gain beyond the 500-sample mark. Thus, we can reasonably conclude that enriched explanation synthesis provides limited added value in full-data scenarios.
Appendix C Cross-Lingual Adaptability
Our method also extends to other languages, particularly character-dense languages in low-resource settings. To assess its cross-linguistic performance, we conducted additional experiments on Japanese and Korean sequence labeling datasets, as summarized in Table 8.
  • Stockmark-NER (Japanese) [^2] https://github.com/stockmarkteam/ner-wikipedia-dataset/: A Wikipedia-style dataset containing eight entity types, including person, organization, and location names.
  • Naver Changwon NER (Korean) [^3] https://github.com/naver/nlp-challenge/tree/master/missions/ner: A social media dataset containing fourteen entity types, including person names, dates, organizations, and locations.
In these experiments, we used GLM4-9B-Chat for data augmentation. The results demonstrate that our method consistently outperforms baseline approaches on both datasets, confirming its effectiveness and adaptability in character-dense, low-resource, multilingual scenarios.
Table 8: Results of sequence labeling tasks on cross-lingual datasets. Bold values indicate the best performance.
Model Stockmark-NER Naver Changwon NER
250 500 1000 250 500 1000
W2NER 68.27 74.84 78.04 66.53 69.17 73.19
CNN Nested NER 69.18 75.79 79.18 68.24 70.83 74.55
DiFiNet 65.83 71.76 76.68 68.67 70.87 72.62
KnowFREE-F (GLM4-9B-Chat) 69.26 76.17 79.97 68.70 70.92 75.66
KnowFREE-FS (GLM4-9B-Chat) 73.82 78.12 79.87 69.98 71.23 75.69
Appendix D Impact of the Number of Heads in Local Attention
@@
Figure 7: Performance variation with different numbers of heads in the local attention module.
To analyze the impact of different numbers of attention heads on model performance, we conducted experiments on four flat NER datasets: Weibo, Youku, Taobao, and Resume, using a sampling size of 1000. The model was trained on the sampled data with extension labels extracted by GLM4-9B-Chat, and the results are presented in Figure 7.
The results suggest that the number of attention heads has a relatively moderate influence on model performance. Notably, increasing the number of heads from 8 to 10 yields the most substantial improvement. Moreover, how feature vectors are distributed across heads proves to be a critical factor. To maintain compatibility as the number of heads increases, we adjusted the feature size to ensure it remains divisible by the number of heads. However, this adjustment did not result in further performance gains and significantly increased computational overhead. Therefore, we adopt ten attention heads in this study as a trade-off between performance and efficiency.
Appendix E Comparison of different data synthesis strategies
To evaluate the effectiveness of the enriched explanation synthesis strategy, we reproduced and compared two LLM-based data synthesis methods designed for sequence labeling tasks: LLM-DA and ProgGen. These methods were used to synthesize data from k-shot samples of the original datasets. For consistency, all synthesized samples were annotated using the KnowFREE model trained on the original data without synthesized samples. We conducted experiments on the Weibo, Youku, Taobao, and Resume datasets, all results are presented in Figure 6. The results show that the performance gains from all data synthesis strategies decrease as the number of samples increases. Notably, in scenarios with $\text{k}\leq 30$ on the Youku and Taobao datasets, both LLM-DA and ProgGen lead to performance degradation compared to models trained without synthesized data. This suggests that the synthesized samples generated by these methods may contain inherent semantic distribution biases, which diminish their effectiveness in enhancing performance in certain low-resource domains.
@@ Appendix E Comparison of different data synthesis strategies
Appendix F Visualization of the Logits with Extension Labels
@@
Appendix G Statistics of Datasets
@@
Appendix H More Experiment Settings
In this section, we describe the additional experimental parameter settings for our method. In the pipeline of label extension annotation, the pre-trained embedding model of $\mathcal{M}$ is set as “text2vec” (Ming, 2022), the Top-$p$ value is set as five, and the threshold $\epsilon$ is set as 1.5. In the training stage, the hidden size $D$, $D^{\prime}$, and $\tilde{D}$ are set as 768, 200, and 200, respectively. The activation function of $\sigma$ and $\sigma^{\ast}$ are defined as Leaky ReLU and GeLU, respectively. In the local multi-head attention module, the window size $\omega$ is set as three and the number of attention heads $\mathcal{K}$ is set as ten. In the sequence labeling model, we distinguish between the learning rate for the PLM and other modules, setting them to 2e-5 and 1e-3, respectively. For the weight $\alpha$ of extension labels, we introduce a dynamic weight calculation mechanism to handle the influence of frequently occurring extension labels (e.g., POS tags). These frequent labels can affect the gradient calculation, leading to reduced attention to target labels. To address this, we calculate the count $C_{i}$ of each extension label and the average count $\hat{C}$ of target labels, and then compute the weight coefficient $\alpha_{i}$ as follows: